OpenStudy (anonymous):

Find the values of x in the following equation: |8x + -2| = 10

OpenStudy (mathstudent55):

To solve an absolute value equation like this one, rewrite it as two simpe equations separated by the word "or." The first equation is just like the original equation, just drop the absolute value signs. The second equation is like the first equation, except you make the number on the right side negative. For example, if you need to solve the equation: |x - 3| = 8, you do this: x - 3 = 8 or x - 3 = -8 x = 11 or x = -5 Now try this with your equation.